During a race, the 80kg sled moves from point A to points B and C on an ice rink. Find: a) The potential energy of the sled at points A, B, C with respect to the zero level of potential energy passes through point A. this level. c) The potential energy of the sled at points A, B and C with respect to the zero level of potential energy passing through the points C d) The change in potential energy of the sled when it shifts from point B to C with respect to the second level zero of potential energy


1
Expert's answer
2021-04-13T06:28:10-0400

a) In this case:


"E_A=mg(H_A-H_A)=0,\\
E_B=mg(H_B-H_A),\\
E_C=mg(H_C-H_A)."

c) In this scenario:


"E_A=mg(H_A-H_C),\\
E_B=mg(H_B-H_C),\\
E_C=mg(H_C-H_C)=0."

d) At point B:


"E_A=mg(H_A-H_B),\\
E_B=mg(H_B-H_B)=0,\\
E_C=mg(H_C-H_B)=0."